A body is propelled through a viscous medium with a speed of 300 cm/sec. On the propelling force being suddenly removed, it comes to rest after covering a distance of 45 metres. What is its relaxation time? How long would it be before its speed falls to 30 cm/sec.


1
Expert's answer
2021-02-28T07:23:24-0500

"t=\frac sv=\frac{45}{3}=15~s,"

"t_1=\frac{v-v_1}{a}=\frac{v-v_1}{\frac vt}=t\cdot \frac{v-v_1}{v}=15\cdot\frac{3-0.3}{3}=13.5~s."